Harry walked to the town at a speed of 4 miles/hr, and came back by bicycle at a speed of 12 miles/hr. The total time for to and fro journey was 5/2 hrs. Find the time taken by him to travel foot.
Answer by stanbon(75887)

Harry walked to the town at a speed of 4 miles/hr, and came back by bicycle at a speed of 12 miles/hr. The total time for to and fro journey was 5/2 hrs. Find the time taken by him to travel foot.
----
Walk DATA:
rate = 4 mph ; distance = x miles ; time = x/4 hrs
------------
Bike DATA:
rate = 12 mph ; distance = x miles; time = 1/12 hrs
-----------------------
Equation:
time + time = 5/2
x/4 + x/12 = 5/2
---
3x + x = 30
4x = 30
x = 7.5 miles (
-----
Walk time = x/4 = (7.5/4)(60 min) = 112.5 min = 1 hr 52.5 min
==================
Cheers,
Stan H.